OpenStudy (anonymous):

Find all solutions of the system of equations. x − 2y = 3 y2 − x2 = 3x + 9

OpenStudy (anonymous):

put x=3+2y in second equation

OpenStudy (anonymous):

what about that x being squared?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

see i am telling you that from first equation, we get x=3+2y put this value of x in second equation

OpenStudy (anonymous):

x. not the x ^2?

OpenStudy (anonymous):

mam second equation is having x^2

OpenStudy (anonymous):

is it y^2-x^2=3x+9 the second equation

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

im writing

OpenStudy (anonymous):

y^2-(2y+3)(2y+3)=3(2y+3)+9?

OpenStudy (anonymous):

y^2-(2y+3)^2=3(2y+3)+9 (y-(2y+3))(y+(2y+3))=6y+9+9 (-y-3)(3y+3)=6y+18

OpenStudy (anonymous):

ok till here????????

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

-(y+3)(3y+3)=6y+18 -(y+3)3(y+1)=3(2y+6) cancel 3 on both sides -(y+3)(y+1)=2y+6

OpenStudy (anonymous):

-(y^2+4y+3)=2y+6 -y^2-4y-3=2y+6 -y^2-4y-2y-3-6=0 -y^2-6y-9=0 -(y^2+6y+9)=0 therefore,y^2+6y+9=0 (y+3)^2=0 so,y=-3,-3

OpenStudy (anonymous):

yes or no?????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so put y=-3 in x=3+2y x=3+2(-3) x=3-6 x=-3 so,(-3,-3) is the solution set

OpenStudy (anonymous):

fine emily.......

OpenStudy (anonymous):

ya got or not? you can ask wherever you have doubt?

OpenStudy (anonymous):

Thank you, sorry i had someone at the door . Thank you for your help :)

OpenStudy (anonymous):

you are always welcome.......any other question?

OpenStudy (anonymous):

nope, thats all for now thank you